OpenStudy (anonymous):

You hold a bag filled with 7 green, 5 red, and blue marbles. What is the probability of selecting a blue marble at random?

OpenStudy (anonymous):

Their is 8 blue marbles.

OpenStudy (ganpat):

total number of marbles = 8 + 7 + 5 = 20.. no of blue marble = 8 so probability of selecting a blue marble = 8/20 = 2/5..

OpenStudy (anonymous):

Thank you!!

OpenStudy (ganpat):

w/c...